what is the kinetic energy of the continent
1
Expert's answer
2019-11-29T09:56:31-0500

The mass of the North American continent, for instance, is estimated to be "2.35095\cdot10^{23}" kg. It moves at the rate of 3.8 cm/year. The kinetic energy is


"KE=mv^2/2=\\
\space\\
=\frac{1}{2}2.35095\cdot10^{23}\cdot\bigg[3.8\cdot\frac{1}{100\cdot365\cdot3600}\bigg]^2=\\\space\\=1706.74\text{ J}."